# Optimal Client Sampling in Federated Learning with Client-Level Heterogeneous Differential Privacy

### v1

@@ 3 Federated Learning with Heterogeneous Group Client-Level DP
Problem Formulation. In this work, we consider an HDPFL setting where each client has its own privacy budget $\epsilon_{i},\ \forall i\in[n]$. The objective is to collaboratively train a global model with satisfactory utility while respecting each client’s privacy preference. To achieve this, our proposed method, GDPFed, partitions all clients into $M$ groups $\mathcal{G}_{1},\mathcal{G}_{2},\dots,\mathcal{G}_{M}$ based on their privacy budgets. Note that the FL problem now is formalized as $\min_{\theta}\sum_{m\in[M]}\omega_{m}\sum_{i\in\mathcal{G}_{m}}f_{m,i}(\theta)$, where $f_{m,i}(\theta)$ is the local learning objective of client $i$ in $\mathcal{G}_{m}$ and $\omega_{m}$ is a reweighting parameter for each group. In each training round $t$ of GDPFed, the server samples a subset of $r_{m}$ clients $\mathcal{S}^{t}_{m}$ from each group $m\in[M]$ where the number of sampled clients $r_{m}$ in group $m$ is determined by the client sampling ratio $q_{m}$ and calculated as $r_{m}=q_{m}|\mathcal{G}_{m}|$. To achieve client-level DP within each group, every local model update in group $m$ is perturbed by adding Gaussian noise drawn from $\mathcal{N}(0,C^{2}\sigma_{m}^{2}/r_{m}\cdot\mathbf{I}^{d})$ after clipping with clipping threshold $C$. Note that the noise multiplier $\sigma_{m}^{2}$ is set to satisfy the minimum privacy budget within each group, denoted by $\epsilon_{m}=\min\{\epsilon_{m,i}\}_{i\in\mathcal{G}_{m}}$, to ensure that clients’ privacy losses are smaller than their budgets. Consequently, selected clients send the perturbed local updates via secure aggregation. One can follow the approach in [9, 25] to implement secure aggregation, and we note that designing a novel secure aggregation protocol is beyond the scope of this paper. The server receives the model update summation from each group and aggregates them with reweighting parameters to refine the global model. This process will repeat for $T$ rounds to ensure that the global model achieves sufficient utility.
Privacy Analysis of GDPFed. We provide per-group privacy guarantees of GDPFed in Theorem 1.
@@
Suppose clients in group $m$ are sampled without replacement with probability $q_{m}$ at each round. For any $\epsilon_{m}<2\log(1/\delta)$ and $\delta\in(0,1)$, GDPFed satisfies $(\epsilon_{m},\delta)$-DP for clients in group $m$ after $T$ rounds if $\sigma^{2}_{m}\geq 7q^{2}_{m}T(\epsilon_{m}+2\log(1/\delta))/\epsilon^{2}_{m}$.
@@
The detailed proof is provided in Appendix C. ∎
@@
In addition to the per-group privacy guarantees provided by GDPFed, we also establish its overall privacy guarantee. To this end, we first present the principle of parallel composition for DP mechanisms, as stated in Lemma 1.
Lemma 1 (Parallel Composition of DP [60, 32]).
Let $\mathcal{M}_{m}:\mathcal{D}_{m}\rightarrow\mathbb{R}^{d}$ be a randomized mechanism that satisfies $(\epsilon_{m},\delta)$-DP, where $\{\mathcal{D}_{m}\}_{m\in[M]}$ are disjoint subsets of the domain $\mathcal{D}$. Then, any randomized function applied to the sequence $\{\mathcal{M}_{m}\}_{m\in[M]}$ satisfies $(\max_{m\in[M]}\epsilon_{m},\delta)$-DP.
That is, if the input domain is partitioned into disjoint subsets independently of the actual data, and each subset is protected using a DP mechanism, the overall privacy guarantee is determined solely by the weakest guarantee (i.e., the highest privacy budget) among the individual mechanisms. Using Lemma 1, we can easily establish the overall privacy guarantee of GDPFed as in Theorem 2.
Theorem 2 (Privacy Guarantee of GDPFed).
If each group $m\in[M]$ in GDPFed selects the noise multiplier $\sigma^{2}_{m}$ satisfies Theorem 1, then after $T$ training rounds, the GDPFed satisfies $\{(\epsilon_{m},\delta)\}_{m\in[M]}$ heterogeneous group-wise DP and $(\max_{m\in[M]}\epsilon_{m},\delta)$-DP.
@@
It is clear that $\min_{i\in[n]}\epsilon_{i}\leq\max_{m\in[M]}\min_{i\in\mathcal{G}_{m}}\epsilon_{i}\leq\max_{i\in[n]}\epsilon_{i}$ where the two equalities hold under the homogeneous DP setting. Compared with DP-FedAvg, which guarantees a $(\min_{i\in[n]}\epsilon_{i},\delta)$-DP for each client in the system, GDPFed achieves a weaker guarantee. Nevertheless, both approaches ensure that any client’s privacy budget is not violated. Importantly, GDPFed relaxes the guarantee for clients with looser privacy requirements, potentially improving the utility of the resulting global model.
@@
Analyzing DP Noise. Building upon the privacy analysis of GDPFed, we now conduct a detailed investigation of the factors that influence the magnitude of the DP noise applied to the model updates, aiming to derive further insights for improving model utility. In GDPFed, we leverage the Gaussian mechanism to impose noise for each group, drawn from the distribution $\mathcal{N}(0,(C^{2}\sigma^{2}_{m}/r_{m})\cdot\mathbf{I}^{d})$, thereby ensuring $(\epsilon_{m},\delta)$-DP. The expected squared $\ell_{2}$-norm of the total noise applied to aggregated model updates (denoted as $\Lambda_{m}$) received by the server is $\Lambda_{m}=d\cdot C^{2}\sigma_{m}^{2}$, for group $m$. Substituting $\sigma_{m}^{2}$ with its lower bound from Theorem 1, we obtain $\Lambda_{m}=7dq_{m}^{2}T(\epsilon_{m}+2\log(1/\delta))C^{2}/\epsilon^{2}_{m}$. We focus on analyzing the influence of two critical parameters, $d$ and $q_{m}$, on the magnitude of DP noise, as other parameters are typically fixed in a given HDPFL system. Specifically, properly adjusting $d$ and $q_{m}$ can effectively reduce the amount of noise under the same privacy guarantee. If model utility is preserved in the process, this can potentially lead to improved overall performance.
@@
a) Reducing $d$. Modern neural network architectures (e.g., ResNet [21]) are typically designed with millions of parameters to ensure strong generalization capability. This results in a large model dimensionality $d$, which in turn significantly increases the magnitude of DP noise. To reduce $d$, existing works consider low-rank decomposition [18, 62, 59] or structured pruning [23, 22]. However, these methods suffer from significant utility loss [25]. Moreover, they alter the model architecture, which poses challenges for model aggregation in FL. A more effective approach is to retain the original architecture while reducing the number of active parameters, a technique known as model sparsification [31] (also known as unstructured pruning). This strategy selectively eliminates a subset of model parameters, which directly reduces DP noise while preserving both the original network architecture and model performance, leveraging the natural redundancy present in DNNs.
@@
b) Adjusting $q_{m}$. Regarding $q_{m}$, directly reducing it leads to a smaller magnitude of DP noise injected into model updates for group $m$. Intuitively, it is desirable to reduce the sampling probability for groups with tighter privacy requirements (i.e., smaller $\epsilon_{m}$), as these groups demand lower privacy loss. In practice, privacy-sensitive clients indeed prefer to participate less frequently in training, reducing their exposure to potential inference attacks [33]. However, this approach degrades global model performance if insufficient clients participate in local training. Assuming a minimum participation ratio $q$ is required (i.e., in expectation, $qn$ clients are selected for local training in each round of GDPFed), clients with larger privacy budgets should participate more frequently, as their model updates contain less noise. Yet, excessive participation frequency also increases DP noise under the same privacy guarantee. Consequently, there exists an optimal set of client sampling ratios that balances these competing factors while satisfying both participation and privacy constraints.
4 Sparsification-Amplified GDPFed with Optimal Client Sampling
In this section, we further improve the model utility of GDPFed by integrating sparsification techniques and deriving the optimal per-group client sampling ratios. This improved version of GDPFed is referred to as GDPFed+, as detailed in Algorithm 1.
@@ 4 Sparsification-Amplified GDPFed with Optimal Client Sampling
GDPFed with Per-group Sparsification. To achieve client-level DP under our attack model, sampled clients in each group add a small amount of noise to the model updates (line 13 in Algorithm 1) and send them to the server via secure aggregation (line 14). The secure aggregation ensures the server only receives the sum of model updates from each group, as well as the summed noise (line 16). Here, per-group DP perturbation is applied over the entire parameter space (i.e., $\mathbb{R}^{d}$) of model updates. In other words, all parameters are subjected to perturbation regardless of their importance. However, prior studies have shown that neural networks typically exhibit substantial parameter redundancy, with many parameters contributing negligibly to the task [31, 19, 20]. Under DP settings, perturbing unimportant parameters introduces redundant noise, unnecessarily degrading model utility.
A practical remedy is model sparsification, which removes unimportant parameters from model updates along with their associated noise. Specifically, a top-$k$ sparsifier, denoted as $\mathrm{Top}_{k}(\cdot)$, is applied to retain only the $k\in[0,d]$ most important parameters. Note that $k=0$ corresponds to eliminating all parameters, while $k=d$ indicates no sparsification. The detailed algorithm of $\mathrm{Top}_{k}(\cdot)$ is provided in Algorithm 3 in Appendix B.3. In this work, we adopt a widely-used and straightforward criterion for identifying important parameters—their absolute magnitude [19, 25, 57, 56]. It should be noted that sparsification must be applied after DP perturbation to preserve the desired $(\epsilon,\delta)$-DP guarantee, as ensured by the post-processing property of DP given as in Lemma 2.
@@
Let $\mathcal{M}$ be a randomized mechanism that satisfies $(\epsilon,\delta)$-DP. Then, for any (possibly randomized) mapping $g$, the composed function $g\circ\mathcal{M}$ also satisfies $(\epsilon,\delta)$-DP.
Technically, in training round $t$, the server applies the $\mathrm{Top}_{k}(\cdot)$ sparsifier to the per-group model updates summation $\bar{\mathbf{y}}^{t}_{m}$ using a group-specific sparsification parameter $k_{m}$, resulting in sparsified updates $\tilde{\mathbf{y}}^{t}_{m}$ (line 17). These sparsified updates are then aggregated with reweighting parameters to refine the global model (line 19).
Bounded Sparsification Error. To reflect varying privacy preferences, it is desirable to assign distinct sparsification parameters (i.e., $k_{1},k_{2},\dots,k_{M}$) to different groups. Intuitively, groups with stricter privacy requirements should be assigned more aggressive sparsification to mitigate the larger DP noise added to their updates. However, $\mathrm{Top}_{k}(\cdot)$ is not without cost since using a smaller $k_{m}$ means that more parameters are removed, which can potentially lead to a non-negligible loss in utility. To formally quantify this relationship, we introduce Lemma 3, which characterizes the approximation error introduced by the $\mathrm{Top}_{k}(\cdot)$ sparsifier.

### v3

@@ IV-A GDPFed
Recall that in this work, we consider an HDPFL setting where each client has its own privacy budget $\epsilon_{i},\ \forall i\in[n]$. The objective is to collaboratively train a global model with satisfactory utility while respecting each client’s privacy preference. To achieve this, our proposed method, GDPFed, partitions all clients into $M$ groups $\mathcal{G}_{1},\mathcal{G}_{2},\dots,\mathcal{G}_{M}$ based on their privacy budgets. Note that the FL problem now is formalized as $\min_{\theta}\sum_{m\in[M]}\omega_{m}\sum_{i\in\mathcal{G}_{m}}f_{m,i}(\theta)$, where $f_{m,i}(\theta)$ is the local learning objective of client $i$ in $\mathcal{G}_{m}$ and $\omega_{m}$ is a reweighting parameter for each group. In each training round $t$ of GDPFed, the server samples a subset of $r_{m}$ clients $\mathcal{S}^{t}_{m}$ from each group $m\in[M]$ where the number of sampled clients $r_{m}$ in group $m$ is determined by the client sampling ratio $q_{m}$ and calculated as $r_{m}=q_{m}|\mathcal{G}_{m}|$. To achieve client-level DP within each group, every local model update in group $m$ is perturbed by adding Gaussian noise drawn from $\mathcal{N}(0,C^{2}\sigma_{m}^{2}/r_{m}\cdot\mathbf{I}^{d})$ after clipping with clipping threshold $C$. Note that the noise multiplier $\sigma_{m}^{2}$ is set to satisfy the minimum privacy budget within each group, denoted by $\epsilon_{m}=\min\{\epsilon_{m,i}\}_{i\in\mathcal{G}_{m}}$, to ensure that clients’ privacy losses are smaller than their budgets. Consequently, selected clients send the perturbed local updates via secure aggregation. One can follow the approach in [8, 25] to implement secure aggregation, and we note that designing a novel secure aggregation protocol is beyond the scope of this paper. The server receives the model update summation from each group and aggregates them with reweighting parameters to refine the global model. This process will repeat for $T$ rounds to ensure that the global model achieves sufficient utility.
IV-B Privacy Analysis of GDPFed
Now we provide the detailed privacy analysis for the proposed GDPFed. Specifically, we first examine the conditions under which each group in GDPFed satisfies the DP requirements. Then, we present privacy guarantees for the entire GDPFed. We provide per-group privacy guarantees of GDPFed in Theorem 1.
@@ IV-B Privacy Analysis of GDPFed
Suppose clients in group $m$ are sampled without replacement with probability $q_{m}$ at each round. For any $\epsilon_{m}<2\log(1/\delta)$ and $\delta\in(0,1)$, GDPFed satisfies $(\epsilon_{m},\delta)$-DP for clients in group $m$ after $T$ rounds if
$\sigma^{2}_{m}\geq 7q^{2}_{m}T(\epsilon_{m}+2\log(1/\delta))/\epsilon^{2}_{m}.$
@@
The detailed proof is provided in Appendix F. ∎
@@
In addition to the per-group privacy guarantees provided by GDPFed, we also establish its overall (system-level) privacy guarantee. To this end, we first present the principle of parallel composition for DP mechanisms, as stated in Lemma 1.
Lemma 1 (Parallel Composition of DP [61]).
Let $\{\mathcal{D}_{m}\}_{m\in[M]}$ be a partition of the input domain $\mathcal{D}$ into disjoint subsets. Suppose each randomized mechanism $\mathcal{M}_{m}:\mathcal{D}_{m}\rightarrow\mathbb{R}^{d}$ satisfies $(\epsilon_{m},\delta)$-DP. Then, any (possibly randomized) function applied to the collection $\{\mathcal{M}_{m}\}_{m\in[M]}$ satisfies $(\max_{m\in[M]}\epsilon_{m},\delta)$-DP.
Intuitively, when the input domain is partitioned into disjoint subsets independently of the actual data, and a DP mechanism protects each subset, the weakest mechanism determines the overall privacy guarantee, that is, the one with the largest privacy budget $\epsilon_{m}$. Using Lemma 1, we can establish the system-level privacy guarantee of GDPFed in Corollary 1.
Corollary 1 (Privacy Guarantee of GDPFed).
Assume that the local datasets of all clients are pairwise disjoint, i.e., no data point is shared between any two clients. If, in GDPFed, each group $m\in[M]$ selects a noise multiplier $\sigma^{2}_{m}$ satisfying the condition in Theorem 1, then after $T$ training rounds, GDPFed provides $\{(\epsilon_{m},\delta)\}_{m\in[M]}$ group-wise DP guarantees, and the entire system satisfies $(\max_{m\in[M]}\epsilon_{m},\delta)$-DP by the parallel composition in Lemma 1.
@@
Intuitively, we observe that $\min_{i\in[n]}\epsilon_{i}\leq\max_{m\in[M]}\min_{i\in\mathcal{G}_{m}}\epsilon_{i}\leq\max_{i\in[n]}\epsilon_{i},$ where both inequalities become equalities under the homogeneous DP setting (i.e., $\epsilon_{i}=\epsilon$ for all clients). Compared with DP-FedAvg, which guarantees $(\min_{i\in[n]}\epsilon_{i},\delta)$-DP for every client, GDPFed yields a slightly weaker overall guarantee. Nevertheless, it flexibly accommodates heterogeneous privacy requirements by allowing groups with looser privacy budgets to use smaller noise multipliers, which in turn can improve the utility of the model. In practice, datasets owned by different groups may contain overlapping data points, which violates the disjointness condition required by Lemma 1. Consider the case where groups $\mathcal{G}_{i}$ and $\mathcal{G}_{j}$ have overlapping datasets and an individual record appears in both groups. In this case, removing this record affects both DP mechanisms $\mathcal{M}_{i}$ and $\mathcal{M}_{j}$, and thus changes their joint output. The ratio between the output distributions on adjacent datasets is given by the product of the likelihood ratios of $\mathcal{M}_{i}$ and $\mathcal{M}_{j}$, which leads to an effective privacy budget that is upper bounded by $\epsilon_{i}+\epsilon_{j}$ (sequential composition), rather than $\max\{\epsilon_{i},\epsilon_{j}\}$ (parallel composition). As a result, the overall privacy guarantee becomes $\left(\max\left\{\epsilon_{i}+\epsilon_{j},\;\max_{m\in[M]\setminus\{i,j\}}\epsilon_{m}\right\},\delta\right)\text{-DP},$ which is weaker than or equal to the ideal disjoint case, since $\max\left\{\epsilon_{i}+\epsilon_{j},\;\max_{m\in[M]\setminus\{i,j\}}\epsilon_{m}\right\}\geq\max_{m\in[M]}\epsilon_{m}.$ Since our goal in this work is to characterize the system-level privacy guarantee of the GDPFed system theoretically, we assume that each client’s dataset is disjoint from all others to ensure that the parallel composition result in Lemma 1 applies. In our empirical evaluation, we also enforce this assumption by explicitly constructing non-overlapping client datasets.
IV-C Analyzing DP Noise
@@
Building upon the privacy analysis of GDPFed, we now conduct a detailed investigation of the factors that influence the magnitude of the DP noise applied to the model updates, aiming to derive further insights for improving model utility. In GDPFed, we leverage the Gaussian mechanism to impose noise for each group, drawn from the distribution $\mathcal{N}(0,(C^{2}\sigma^{2}_{m}/r_{m})\cdot~\mathbf{I}^{d})$, thereby ensuring $(\epsilon_{m},\delta)$-DP. The expected squared $\ell_{2}$-norm of the total noise applied to aggregated model updates (denoted as $\Lambda_{m}$) received by the server is $\Lambda_{m}=d\cdot C^{2}\sigma_{m}^{2}$, for group $m$. Substituting $\sigma_{m}^{2}$ with its lower bound from Theorem 1, we obtain $\Lambda_{m}=7dq_{m}^{2}T(\epsilon_{m}+2\log(1/\delta))C^{2}/\epsilon^{2}_{m}$. We focus on analyzing the influence of two critical parameters, $d$ and $q_{m}$, on the magnitude of DP noise, as other parameters are typically fixed in a given HDPFL system. Specifically, properly adjusting $d$ and $q_{m}$ can effectively reduce the amount of noise under the same privacy guarantee. If model utility is preserved in the process, this can potentially lead to improved overall performance.
@@
a) reducing $d$. Modern neural network architectures (e.g., ResNet [21]) are typically designed with millions of parameters to ensure strong generalization capability. This results in a large model dimensionality $d$, which in turn significantly increases the magnitude of DP noise. To reduce $d$, existing works consider low-rank decomposition [17, 65, 60] or structured pruning [22, 23]. However, these methods suffer from significant utility loss [25]. Moreover, they alter the model architecture, which poses challenges for model aggregation in FL. A more effective approach is to retain the original architecture while reducing the number of active parameters, a technique known as model sparsification [31] (i.e., unstructured pruning). This strategy selectively eliminates a subset of model parameters, which directly reduces DP noise while preserving both the original network architecture and model performance, leveraging the natural redundancy present in deep neural networks.
@@
b) Adjusting $q_{m}$. Regarding $q_{m}$, directly reducing it leads to a smaller magnitude of DP noise injected into model updates for group $m$. Intuitively, it is desirable to reduce the sampling probability for groups with tighter privacy requirements (i.e., smaller $\epsilon_{m}$), as these groups demand lower privacy loss. In practice, privacy-sensitive clients indeed prefer to participate less frequently in training, reducing their exposure to potential inference attacks [33]. However, this approach degrades global model performance if insufficient clients participate in local training. Assuming a minimum global participation ratio $q$ is required (i.e., in expectation, $qn$ clients are selected for local training in each round of GDPFed), clients with larger privacy budgets should participate more frequently, as their model updates contain less noise. Yet, excessive participation frequency also increases DP noise under the same privacy guarantee. Consequently, there exists an optimal set of client sampling ratios that balances these competing factors while satisfying both participation and privacy constraints.
@@
Algorithm 2 Top-$k$ Sparsifier $\mathrm{Top}_{k}(\cdot)$
0:  Vector $x\in\mathbb{R}^{d}$, top-$k$ parameter $k\in[0,d]$
0:  Binary mask $\mathbf{mk}\in\{0,1\}^{d}$
1:  Initialization: Initialize $\mathbf{mk}\leftarrow\mathbf{0}^{d}$
2:  Compute absolute values: $a_{j}\leftarrow|x_{j}|$ for all $j\in[1,d]$
4:  for $j=1$ to $k$ do
5:   $\mathbf{mk}]_{\pi(j)}\leftarrow 1$
6:  end for
7:  return $\mathbf{mk}\odot x$
V Sparsification-Amplified GDPFed with Optimal Client Sampling
In this section, we introduce an enhanced variant of GDPFed, termed GDPFed+, which improves model utility by incorporating sparsification techniques that reduce the model dimension $d$ and by deriving optimal client sampling ratios for each group. The algorithm of GDPFed+ is detailed in Algorithm 1.
V-A GDPFed with Per-group Sparsification
To achieve client-level DP under our attack model, sampled clients in each group add a small amount of noise to the model updates (line 13 in Algorithm 1) and send them to the server via secure aggregation (line 14). The secure aggregation ensures the server only receives the sum of model updates from each group, as well as the summed noise (line 16). Here, per-group DP perturbation is applied over the entire parameter space (i.e., $\mathbb{R}^{d}$) of model updates. In other words, all parameters are subjected to perturbation regardless of their importance. However, prior studies have shown that neural networks typically exhibit substantial parameter redundancy, with many parameters contributing negligibly to the task [31, 19, 20]. Under DP settings, perturbing unimportant parameters introduces redundant noise, unnecessarily degrading utility.
A practical remedy is model sparsification, which removes unimportant parameters from model updates along with their associated noise. Specifically, a top-$k$ sparsifier, denoted as $\mathrm{Top}_{k}(\cdot)$, is applied to retain only the $k\in[0,d]$ most important parameters. Note that $k=0$ corresponds to eliminating all parameters, while $k=d$ indicates no sparsification. The detailed algorithm of $\mathrm{Top}_{k}(\cdot)$ is provided in Algorithm 2. In this work, we adopt a widely-used and straightforward criterion for identifying important parameters—their absolute magnitude [19, 25, 58]. It should be noted that sparsification must be applied after DP perturbation to preserve the desired $(\epsilon,\delta)$-DP guarantee, as ensured by the post-processing property of DP given as in Lemma 2.
@@ V-A GDPFed with Per-group Sparsification
Let $\mathcal{M}$ be a randomized mechanism that satisfies $(\epsilon,\delta)$-DP. Then, for any mapping $g$, the composed function $g\circ\mathcal{M}$ also satisfies $(\epsilon,\delta)$-DP.
Technically, in training round $t$, the server applies the $\mathrm{Top}_{k}(\cdot)$ sparsifier to the per-group model updates summation $\bar{\mathbf{y}}^{t}_{m}$ using a group-specific sparsification parameter $k_{m}$, resulting in sparsified updates $\tilde{\mathbf{y}}^{t}_{m}$ (line 17). These sparsified updates are then aggregated with reweighting parameters to refine the global model (line 19). In the following, we theoretically analyze the error introduced by sparsification.
V-B Bounded Sparsification Error
To reflect varying privacy preferences, it is desirable to assign distinct sparsification parameters (i.e., $k_{1},k_{2},\dots,k_{M}$) to different groups. Intuitively, groups with stricter privacy requirements should be assigned more aggressive sparsification to mitigate the larger DP noise added to their updates. However, $\mathrm{Top}_{k}(\cdot)$ is not without cost since using a smaller $k_{m}$ means that more parameters are removed, which can potentially lead to a non-negligible loss in utility. To formally quantify this relationship, we introduce Lemma 3, which characterizes the approximation error introduced by the $\mathrm{Top}_{k}(\cdot)$ sparsifier.